OpenStudy (anonymous):

at 3:45 siri was going up an elevator from the 2nd floor. At 3:48 she arrived on the 42 floor. What was the average time per floor?

OpenStudy (anonymous):

can i get any help each problem i put on here. i never receive an answer

jimthompson5910 (jim_thompson5910):

She changes 42 - 2 = 40 floors in 48-45 = 3 minutes So divide that 3 minutes (or 60*3 = 180 seconds) up into 40 equal chunks to get 180/40 = 4.5 So she spends 4.5 seconds on each floor

OpenStudy (anonymous):

at 3am a measurement was taken to be 50 miles. At 6am the same measurement was 35 miles. What is that average rateof change per hour? How many total hours go by before the distance register 0 miles?

jimthompson5910 (jim_thompson5910):

You lose 50-35 = 15 miles for every 6-3 = 3 hours So the rate of change is 15/3 = 5 hours and this rate of change is negative (to show that you're losing miles as time goes on)

jimthompson5910 (jim_thompson5910):

oops I meant to write that the rate of change is -5 miles/hr (basically you lose 5 miles every hour)

OpenStudy (anonymous):

is there 2 answers to that q or the only answer is 5 miles/hr

jimthompson5910 (jim_thompson5910):

Well you use that first part to answer the second part

jimthompson5910 (jim_thompson5910):

Since you know that you lose 5 mi per hr, this means that you lose 5x miles in x hrs So 50 - 5x = 0 ----> x = 10 Which means that in 10 hours, the distance will be 0

OpenStudy (anonymous):

determine if the foolfollowing are parallel, perpendicular, or neither (a) y=5x+6 y=-5x+6

OpenStudy (anonymous):

i meant following

jimthompson5910 (jim_thompson5910):

What are the slopes of each line?

OpenStudy (anonymous):

do not know how to do this at all. do i have to get the slope first

jimthompson5910 (jim_thompson5910):

yes, do you know how to find the slope?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

The slope is simply the number in front of the x term

jimthompson5910 (jim_thompson5910):

Since the equation is in y = mx+b form

OpenStudy (anonymous):

the 5 is in front of the x for the first one, how do i determine if its perpendicular, parallel or neither

jimthompson5910 (jim_thompson5910):

So the slope of y = 5x+6 is 5 The slope of y = -5x+6 is -5 Since these two slopes are NOT equal, this means the two lines are NOT parallel.

jimthompson5910 (jim_thompson5910):

Also, -5 is not the negative reciprocal of 5. So the two lines are NOT perpendicular.

OpenStudy (anonymous):

the following is the same q. y=x+4, y=-x-3..........y=1/2x+8,y=2x...........y=7x-2,y=-1/7x+8............y=4,y=-3